The time that it takes for the next train to come follows a Uniform distribution with f(x) =1/25 where x goes between 6 and 31 minutes. Round answers to 4 decimal places when possible. This is a Correct distribution. It is a Correct distribution. The mean of this distribution is 18.50 Correct The standard deviation is Incorrect Find the probability that the time will be at most 28 minutes. Incorrect Find the probability that the time will be between 12 and 24 minutes. Incorrect Find the 12th percentile. Incorrect Find the probability that the time is more than 24 minutes given (or knowing that) it is at least 14 minutes. Incorrect
Answer #1

___fea) = __ _x. luni ( 6;31) ६(५) = 3 4 | b) stx) = (2) v(a) = EN) - Ex)२ s, ExR) = ( ३ कर । 25 __ 23731 3g५.33 .751 २) = [7-Q) = Pex ८२) - ० . २ - न . ..२ - त्र ___6 - 6 25. - 0 - 2 ___ + 9 3) ____ > २५ x 21५) c da ) 95 + __27 32 . . । 48